40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 1186     Entry time: Wed Dec 1 13:34:04 2010
Author: Zach 
Type: Laser 
Category: GYRO 
Subject: current gyro noise budget 

 Attached is a current gyro noise budget. It has not yet been fully automated, but Koji wanted to have something to look at this afternoon so I made this one by hand. One very nice thing to note here is that the current gyro sensitivity is clearly limited by the dark noise of the PDA255 REFL diodes above ~20 Hz, which we suspected but until now haven't substantiated.

A couple notes:

  • The PDA255 dark noise was only plotted as it affects the primary loop, as the second loop appears to have ~5x higher optical gain [V/Hz]
  • Shot noise and demod noise do not currently appear as the current REFL noise is dominated by the PDA255 dark noise, as shown here.

gyro_nb_12_1_10.png

ELOG V3.1.3-